OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (−2, 1)?

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) &({\color{red}{ -2}}\quad ,&{\color{blue}{ 1}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

Get your slope first \( \huge slope(m) = \frac{y_2- y_1}{x_2-x_1}\) \( \huge slope(m) = \frac{1- 6}{-2-3}\) What does the slope equal? After you get the slope we can use the point slope form to solve. Let us know when you have the slope. What does the slope equal?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

y+6=1(x+3) is this correct?

OpenStudy (anonymous):

Point slope form = \( \huge y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

m = 1

OpenStudy (anonymous):

Do you have \( \huge y - 6 = 1(x - 3)\) ?? If not try using the other point